# Luminous Fast Blue Optical Transients as very massive star core-collapse events

### v1

@@ Abstract
Results. The formation rate of black holes with masses greater than 30–40 M⊙ is found to be similar to the observed LFBOT rate. The stars producing these black holes are biased to low metallicity (Z0.3 Z⊙), are H and He-poor and have dense circumstellar media. However, some LFBOTs have host galaxies with higher metallicities than predicted, and others have denser local environments (plausibly due to late stage mass loss not captured in the models). We find that long-lived emission from an accretion disc (as implicated in the prototypical LFBOT AT 2018cow) can plausibly be produced in these events.
Conclusions. We conclude that (very) massive star core-collapse, based on the above criteria, is a plausible explanation for LFBOTs. The preferred progenitors for LFBOTs in this scenario overlap with those predicted to produce super-kilonovae. We therefore suggest that LFBOTs are promising targets to search for super-kilonovae, and that they may contribute non-negligibly to the r-process enrichment of galaxies.
@@
Luminous Fast Blue Optical Transients (LFBOTs), or Cow-like transients after the prototypical AT 2018cow (Prentice et al., 2018; Perley et al., 2019), are rapidly evolving transients with peak optical-ultraviolet luminosities rivalling superluminous supernovae (SLSNe). The early optical-UV spectra are well described by a hot, featureless black body (T20000 K), lacking H and He lines initially. They have decay rates of 0.3 magnitudes/day, ruling out a significant contribution from 56Ni-powered emission. They are also accompanied by bright X-ray (and at late times, radio) emission. The late-time radio emission is attributed to self-absorbed synchrotron emission from an expanding blastwave in a dense circumstellar medium (CSM).
@@
A modest sample of LFBOTs has now been identified in addition to AT 2018cow: ZTF18abvkwla (AT2018lug, Ho et al., 2020), CSS161010 (Coppejans et al., 2020; Gutiérrez et al., 2024), ZTF20acigmel (AT2020xnd, Perley et al., 2021; Bright et al., 2022), AT2020mrf (Yao et al., 2022), AT2022tsd (Matthews et al., 2023), AT2023fhn (Chrimes et al., 2024b, a) and AT2024wpp (Pursiainen et al., 2025; LeBaron et al., 2025; Nayana et al., 2025), as well as the slower-evolving and radio-faint, but otherwise similar, AT2024puz (Somalwar et al., 2025) and several other candidates which lack extensive multi-wavelength follow-up. Some transients display very blue colours, evolve rapidly at early times and produce bright X-ray emission, similar to LFBOTs, but are followed by broad lined type Ic SNe (e.g. the Einstein Probe-discovered transients EP2404014A and EP250108A, van Dalen et al., 2025; Eyles-Ferris et al., 2025; Rastinejad et al., 2025). However, these events likely do not share the same progenitors as LFBOTs, given that they successfully launched supernovae, and we henceforth focus solely on the SN-less, Cow-like LFBOTs, whose origins remain unknown. The volumetric rate of LFBOTs is estimated to be 0.1% of the core-collapse SN rate (Ho et al., 2023b).
@@
Two categories of model to explain LFBOTs have emerged as most likely (e.g. Perley et al., 2019). The first invokes tidal disruption events (TDEs) of H and He poor, compact stars (e.g. a white dwarfs) by intermediate mass black holes (IMBHs, Kuin et al., 2019). This model is challenged primarily due to the dense CSM observed in LFBOTs. Although such a environment is possible in TDEs involving stars disrupted by supermassive black holes, it is less clear whether white dwarf plus IMBH TDEs could produce similar circumstellar density profiles (Margutti et al., 2019; Linial & Quataert, 2024). The second category of proposals are broadly themed around stellar mass black hole accretion, either in mergers/collisions (Metzger, 2022; Grichener, 2025; Tsuna & Lu, 2025) or core-collapse events (e.g. Kashiyama & Quataert, 2015; Margutti et al., 2019; Gottlieb et al., 2022).
@@
Sufficiently massive stars are expected to directly form black holes without launching a supernova. The landscape of explodability is thought to be complex, with neutron star and some black hole-forming events capable of launching successful supernovae, and other black hole-forming core-collapses resulting in direct black hole formation without any significant ejecta (e.g. Heger et al., 2003; Ertl et al., 2020). Observational evidence for ‘quiet’ black hole formation comes from a lack of massive stars M⊙ in pre-supernovae imaging, and the disappearance of supergiant stars without evidence for a supernova (Smartt et al. 2009; Williams et al. 2014; Reynolds et al. 2015; Adams et al. 2017; Beasor et al. 2024). Some LFBOT models invoke such a scenario, where a massive star collapses to a black hole without a successful supernova (a ‘failed supernova’), followed by accretion of the remaining (‘leftover’) material onto the nascent BH (e.g. Fernández et al., 2018; Quataert et al., 2019). The failed supernova hypothesis resembles early collapsar gamma-ray burst (GRB) models (e.g. Woosley, 1993; MacFadyen & Woosley, 1999; Fryer et al., 1999), which also feature accreting black holes born in core-collapse, but without a successful (or a weak) supernova. In LFBOTs there is no evidence for a supernova, but unlike in collapsar GRBs, no gamma-ray emission has been observed, and the outflows are only mildly relativistic (e.g. Coppejans et al., 2020).
@@
While some fast evolving optical transients can be explained solely with circumstellar interactions (Pellegrino et al., 2022), in LFBOTs there is evidence for central engine activity. AT 2018cow showed short timescale variability in its light-curve (Ho et al., 2019; Margutti et al., 2019; Pasham et al., 2021), as well as UV and X-ray emission lasting at least several years post-explosion (Sun et al., 2022, 2023; Inkenhaag et al., 2023; Chen et al., 2023; Migliori et al., 2024). Magnetar central engines have been proposed (Fang et al., 2019; Mohan et al., 2020; Liu et al., 2022; Li et al., 2024), but struggle to explain both the early-time emission and late-time optical/UV plateau (Chen et al., 2023). Assuming instead that LFBOTs are black hole (BH) powered, estimates for the BH mass in such a scenario have been derived from quasi period oscillations (Pasham et al., 2021) and accretion disc modelling (Inkenhaag et al., 2023; Migliori et al., 2024; Cao et al., 2024) yielding masses which span the upper end of the stellar mass BH distribution into the intermediate BH mass regime. The observation of late-time, short duration giant optical flares from AT 2022tsd (Ho et al., 2023a) may also be consistent with a highly variable accretion rate, either around the black hole formed in the event or a companion black hole (Lazzati et al., 2024). Similar flaring behaviour has not been found in AT20224wpp (Ofek et al., 2025).
@@
Given the high BH mass estimates in LFBOTs, and their discovery thus far in non-nuclear regions of low-metallicity, star-forming galaxies, the primary aim of this paper is determine whether massive BH formation in failed supernovae is consistent with the volumetric rate and environments of LFBOTs. In Section 2 we compare population synthesis predictions for the rate of BH formation as a function of mass with the observed LFBOT rate. We then investigate in Section 3 whether the metallicity bias associated with the formation of the most massive stellar mass BHs is consistent with LFBOT environments. We explore in Section 4 whether the selected stellar models can plausibly reproduce the key characteristics of LFBOT emission, with a discussion and conclusions following in Sections 5 and 6.
@@ 2 Black hole formation and LFBOT event rates
For the population synthesis we use the Binary Population and Spectral Synthesis (bpass) models (v2.2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018). These are a publicly available grid of pre-calculated binary stellar evolution models, in which the primary evolution is followed in detail, while the secondary is evolved with the rapid evolution prescriptions of Hurley et al. (2002). Models are provided at 13 metallicities and at each metallicity, each stellar model is weighted according to observed binary parameter distributions (Moe & Di Stefano, 2017). We adopt the model set with a Kroupa (2001) broken power-law initial mass function (IMF), with a minimum mass of 0.1 M⊙ and a maximum of 300 M⊙, a break mass of 0.5 M⊙ and a slope of () below (above) the break. At each metallicity , the weighting of each model corresponds to the number of systems resembling that model expected in a stellar population of M⊙. For full details about the binary stellar models and the code, we refer the reader to the bpass v2 release paper Eldridge et al. (2017).
@@
For the metallicity-dependent cosmic star formation rate history, CSFH, we adopt the model of Langer & Norman (2006). We seed the population synthesis models according to this CSFH such that the volumetric birth rate of each system (bpass model) at each redshift and metallicity is proportional to the product of the star formation rate density SFRD and the model weighting (defined by the IMF and binary parameter distributions as described above).
@@
We adopt the standard bpass assumptions that core-collapse occurs when, at the end of the model, the total mass exceeds 1.5 M⊙, the CO core mass exceeds 1.38 M⊙ and the ONe core mass is non-zero. Remnant masses are a pre-calculated output of the models. They are determined by injecting 1051 erg of energy, calculating the mass in the outer layers which can be lifted to infinity by this energy injection, and taking the remaining mass as the remnant mass (for full details see Eldridge & Tout, 2004; Eldridge et al., 2017). We classify all core-collapse remnants with M3 M⊙ as neutron stars, and those heavier as black holes. We assume that all NS-forming events produce successful supernovae, and that BH-forming events do not. Pair-instability supernovae (PISNe) are deemed to occur if the final carbon-oxygen core mass exceeds 60 M⊙ and the final helium core mass is 133 M⊙. In this case no remnant is left behind (Heger & Woosley, 2002). Pulsational pair instability supernovae (PPISNe) are accounted for by adopting the remnant mass prescription of Farmer et al. (2019) for final carbon-oxygen masses between 38 and 60 M⊙, following Briel et al. (2023). The true landscape of explodability and remnant formation is more complex, but to first order our approximations are expected to be reasonable (e.g. Smartt et al., 2009; Smartt, 2009; Fryer et al., 2012; Ertl et al., 2020; Sukhbold & Adams, 2020; Patton & Sukhbold, 2020; Patton et al., 2022; Laplace et al., 2025; Ugolini et al., 2025; Steinwandel & Goldberg, 2025; The LIGO Scientific Collaboration et al., 2025).
@@
The remnant formation time for each stellar model is the seed time, plus the age of the model at core-collapse (although for these short-lived stars, this is negligible compared with cosmological timescales). In this way, we construct a metallicity-weighted supernova rate and BH formation rate as a function of redshift. Since all (confirmed) LFBOTs thus far have redshifts , we compare the mean rate of supernovae and BH-forming events over this redshift range, for different values of MBH,min. Specifically, we calculate the mean ratio of the supernova rate and the formation rate of BHs with mass MBH,min or greater. As the form of CSFH is uncertain, we repeat the process with two extreme variations of the CSFH. These are defined by shifting the Langer & Norman (2006) model by 0.2 dex in 12log(O/H), which captures the approximate envelope of possible CSFHs found by Chruslinska & Nelemans (2019).
@@
The results are shown in Figure 1. Using the fiducial CSFH we find that the birth rate of BHs at with masses 38 M⊙ is 0.1% of the CCSN rate, matching LFBOTs. For a CSFH of the same form but peaking at a metallicity 0.2 dex lower in 12log(O/H), we find a higher BH mass threshold of 41 M⊙ is required, because there are more high-mass BHs being formed overall with respect to CCSNe. The reverse is true for a CSFH which peaks 0.2 dex higher, yielding a threshold of 33 M⊙. Therefore, varying the CSFH changes the absolute number of BHs being born (although not the general shape of the BH mass distribution, van Son et al., 2023), and the number formed with respect to CCSNe.
@@ 3 Metallicity dependence and LFBOT host galaxies
| AT 2018cow | 0.470.01 | Lyman et al. (2020)⋆ |
| ZTF18abvkwla | 0.37 | Ho et al. (2020) |
| CSS161010 | 0.15 | Gutiérrez et al. (2024) |
| ZTF20acigmel | 0.15 | Perley et al. (2021)† |
| AT2020mrf | 0.27 | Yao et al. (2022) |
| AT2023fhn | 0.080.02 | Chrimes et al. (2024a) |
| AT2024wpp | 0.53 | Pursiainen et al. (2025)† |
⋆ - IFU metallicity measurement, resolves the host giving a local (few hundred parsec scale) value. † - Metallicity derived from the host’s total stellar mass using equation 3 of Tremonti et al. (2004). Uncertainties are solely propagated from uncertainties on the galaxy stellar mass, and do not include scatter in the relation.
@@
Given a black hole mass cut MBH,min which yields a BH formation rate matching the observed LFBOT rate, for an assumed CSFH, we can next ask: what is the (CSFH weighted) metallicity distribution of the progenitors of these black holes, and how does this compare with the host galaxies of LFBOTs? We collate the available information in the literature on LFBOT host metallicities in Table 1. Four of the seven have host-averaged metallicities derived from spectral energy distribution (SED) fitting. In one case (AT 2018cow) there is integral field unit data available, enabling a more local (resolution of a few hundred parsec, Lyman et al., 2020) measurement of the metallicity. For the remaining two, no direct measurement has been published, so we obtain an approximate value through the galaxy mass-metallicity relation (Tremonti et al., 2004). We adopt Z by mass fraction for Solar metallicity, but note that this specific choice has no impact on the comparisons or analysis since both the observed and synthetic metallicities are scaled by the same factor.
@@
In Figure 2 we compare the number of black holes more massive than MBH,min born per M⊙ of star formation, weighted by the mean metallicity distribution for each of the three CSFHs adopted, with the measured distribution of LFBOT host or environmental metallicities. There is a strong preference for BHs with masses greater than 38 M⊙ to be born in stellar populations with metallicity less than 0.3 Z⊙. Below this threshold, the distributions are relatively constant: this reflects that the fact that (i) we are tailoring our mass cut to yield a BH formation rate that is 0.1% of the CCSN rate for each CSFH, and (ii) that the CCSN rate itself is not substantially varying with the CSFH variations. This is because NS formation efficiency, unlike BH formation efficiency, does not sharply decline at higher metallicities (e.g. van Son et al., 2025).
@@
It is immediately apparent that LFBOT environments sample the ZZ⊙ region. The formation of BHs with MMBH,min across this range instead samples metallicities ZZ⊙. A KS-test between the LFBOT host values and the fiducial M⊙ distribution returns , formally rejecting the null hypothesis that the observed host values are consistent with the synthetic distribution. However, we note that the host metallicities are far from uniformly determined - with methods ranging from resolved IFU spectroscopy, probing the local environment within the host, to host-integrated metallicites to the mass-metallicity relation. This likely introduces errors/scatter which has not been accounted for, and host-integrated measurements will naturally be higher than the lowest metallicity environments within each host. There may also be pockets of low metallicity gas within and around star-forming galaxies, driving low-Z star-formation (e.g. Michałowski et al., 2015; Metha & Trenti, 2023). Nevertheless, the current information suggesting that LFBOTs can occur in environments with metallicities above the strong cut-off predicted is a challenge to the model. This is most notable for AT2018cow, whose metallicity measurement is integrated across the immediate (few hundred pc) environment (Lyman et al., 2020), but even in this case variations on smaller scales are possible.

### v2

@@ Abstract
Results. The formation rate of black holes with masses greater than 30–40 M⊙ is found to be similar to the observed LFBOT rate. The stars producing these black holes are biased to low metallicity (Z0.3 Z⊙), are H and He-poor and have dense circumstellar media. However, some LFBOTs have host galaxies with higher metallicities than predicted, and they typically have denser local environments (plausibly due to late stage mass loss not captured in the models). We find that long-lived emission from an accretion disc (as implicated in the prototypical LFBOT AT 2018cow) can only be produce in these events under maximal disc mass and angular momentum conditions.
Conclusions. We conclude that (very) massive star core-collapse is a plausible explanation for at least some LFBOTs, but faces challenges. The preferred progenitors for LFBOTs in the failed supernova interpretation overlap with those predicted to produce super-kilonovae. We therefore suggest that LFBOTs are promising targets to search for super-kilonovae, and that they may contribute non-negligibly to the r-process enrichment of galaxies.
@@
Luminous Fast Blue Optical Transients (LFBOTs), or Cow-like transients after the prototypical AT 2018cow (2018ApJ...865L...3P; 2019MNRAS.484.1031P), are rapidly evolving transients with peak optical-ultraviolet luminosities rivalling superluminous supernovae (SLSNe). The early optical-UV spectra are well described by a hot, featureless blackbody (T20000 K), lacking H and He lines initially. They have decay rates of 0.3 magnitudes/day, ruling out a significant contribution from 56Ni-powered emission. They are also accompanied by bright X-ray (and at late times, radio) emission. The late-time radio emission is attributed to self-absorbed synchrotron emission from an expanding blastwave in a dense circumstellar medium (CSM).
@@
A modest sample of LFBOTs has now been identified in addition to AT 2018cow: ZTF18abvkwla (AT2018lug, 2020ApJ...895...49H), CSS161010 (2020ApJ...895L..23C; 2024ApJ...977..162G), ZTF20acigmel (AT2020xnd, 2021MNRAS.508.5138P; 2022ApJ...926..112B), AT2020mrf (2022ApJ...934..104Y), AT2022tsd (2023RNAAS...7..126M), AT2023fhn (2024MNRAS.527L..47C; 2024AandA...691A.329C) and AT2024wpp (2025MNRAS.537.3298P; 2025ApJ...993L...6N; 2025arXiv250900952N), as well as the slower-evolving and radio-faint, but otherwise similar, AT2024puz (2025ApJ...995..228S) and several other candidates which lack extensive multi-wavelength follow-up. Some transients display very blue colours, evolve rapidly at early times and produce bright X-ray emission, similar to LFBOTs, but are followed by broad lined type Ic SNe (e.g. the Einstein Probe-discovered transients EP2404014A and EP250108A, 2025ApJ...982L..47V; 2025ApJ...988L..14E; 2025ApJ...988L..13R). However, these events likely do not share the same progenitors as LFBOTs, given that they successfully launched supernovae, and we henceforth focus solely on the SN-less, Cow-like LFBOTs, whose origins remain unknown. The volumetric rate of LFBOTs is estimated to be 0.1% of the core-collapse SN rate (2023ApJ...949..120H).
@@
Two categories of model to explain LFBOTs have emerged as most likely (e.g. 2019MNRAS.484.1031P). The first invokes tidal disruption events (TDEs) of H and He poor, compact stars (e.g. a white dwarfs) by intermediate mass black holes (IMBHs, 2019MNRAS.487.2505K). This model is challenged primarily due to the dense CSM observed in LFBOTs. Although such a environment is possible in TDEs involving stars disrupted by supermassive black holes, it is less clear whether white dwarf plus IMBH TDEs could produce similar circumstellar density profiles (2019ApJ...872...18M; 2024ApJ...974...67L). The second category of proposals are broadly themed around stellar mass black hole accretion, either in mergers/collisions (2022ApJ...932...84M; 2025Ap&SS.370...11G; 2025ApJ...986...84T; 2025arXiv251009745K) or core-collapse events (e.g. 2015MNRAS.451.2656K; 2019ApJ...872...18M; 2019MNRAS.485L..83Q; 2022MNRAS.513.3810G).
@@
At least some sufficiently massive stars are expected to directly form black holes without launching a supernova. The landscape of explodability is thought to be complex, with neutron star and some black hole-forming events capable of launching successful supernovae, and other black hole-forming core-collapses resulting in direct black hole formation without any significant ejecta (e.g. 2003ApJ...591..288H; 2020ApJ...890...51E). Observational evidence for ‘quiet’ black hole formation comes from a lack of massive stars M⊙ in pre-supernovae imaging, and the disappearance of supergiant stars without evidence for a supernova (2009MNRAS.395.1409S; 2014ApJ...791..105W; 2015MNRAS.453.2885R; 2017MNRAS.468.4968A; 2017MNRAS.469.1445A; 2024ApJ...964..171B).
Some LFBOT models invoke such a scenario, where a massive star collapses to a black hole without a successful supernova (a ‘failed supernova’), followed by the accretion of some fraction of the remaining (‘leftover’) material onto the nascent BH (e.g. 2014ApJ...781..119P; 2015MNRAS.451.2656K; 2018MNRAS.476.2366F; 2019MNRAS.485L..83Q). The failed supernova hypothesis resembles early collapsar gamma-ray burst (GRB) models (e.g. 1993ApJ...405..273W; 1999ApJ...524..262M; 1999ApJ...526..152F), which also feature accreting black holes born in core-collapse, but without a successful (or a weak) supernova. In LFBOTs there is no evidence for a supernova, but unlike in collapsar GRBs, no gamma-ray emission has been observed, and the outflows are only mildly relativistic (e.g. 2020ApJ...895L..23C).
@@
While some fast evolving optical transients can be explained solely with circumstellar interactions (2022ApJ...926..125P), in LFBOTs there is evidence for central engine activity. AT 2018cow showed short timescale variability in its light-curve (2019ApJ...871...73H; 2019ApJ...872...18M; 2022NatAs...6..249P), as well as UV and X-ray emission lasting at least several years post-explosion (2022MNRAS.512L..66S; 2023MNRAS.519.3785S; 2023MNRAS.525.4042I; 2023ApJ...955...43C; 2024ApJ...963L..24M; 2025MNRAS.544L.108I). Magnetar central engines have been proposed (2019ApJ...878...34F; 2020ApJ...888L..24M; 2022ApJ...935L..34L; 2024ApJ...963L..13L), but struggle to explain both the early-time emission and late-time optical/UV plateau (2023ApJ...955...43C). Assuming instead that LFBOTs are black hole (BH) powered, estimates for the BH mass in such a scenario have been derived from quasi period oscillations (2022NatAs...6..249P) and accretion disc modelling (2023MNRAS.525.4042I; 2024ApJ...963L..24M; 2024A&A...691A.228C; 2025MNRAS.544L.108I) yielding masses which span the upper end of the stellar mass BH distribution into the intermediate BH mass regime. The observation of late-time, short duration giant optical flares from AT 2022tsd (2023Natur.623..927H) may also be consistent with a highly variable accretion rate, either around the black hole formed in the event or a companion black hole (2024ApJ...972L..17L). Similar flaring behaviour has not been found in AT20224wpp (2025ApJ...993...76O).
@@
Black hole mass estimates in LFBOTs are high, and they have thus far been discovered in non-nuclear regions of low-metallicity, star-forming galaxies. The aim of this paper is to explore other observational consequences - for example in terms of volumetric rates and the metallicities of their hosts - if LFBOTs are indeed powered by massive BH formation in failed supernovae. In Section 2 we compare population synthesis predictions for the rate of BH formation as a function of mass with the observed LFBOT rate. We then investigate in Section 3 whether the metallicity bias associated with the formation of the most massive stellar mass BHs is consistent with LFBOT environments. We explore in Section 4 whether the selected stellar models can plausibly reproduce the key characteristics of LFBOT emission, with a discussion and conclusions following in Sections 5 and 6.
@@ 2 Black hole formation and LFBOT event rates
For the population synthesis we use the Binary Population and Spectral Synthesis (bpass) models (v2.2.1, 2017PASA...34...58E; 2018MNRAS.479...75S). These are a publicly available grid of pre-calculated binary stellar evolution models, in which the primary evolution is followed in detail, while the secondary is evolved with the rapid evolution prescriptions of 2002MNRAS.329..897H. Models are provided at 13 metallicities and at each metallicity, each stellar model is weighted according to observed binary parameter distributions (2017ApJS..230...15M). We adopt the model set with a 2001MNRAS.322..231K broken power-law initial mass function (IMF), with a minimum mass of 0.1 M⊙ and a maximum of 300 M⊙, a break mass of 0.5 M⊙ and a slope of () below (above) the break. At each metallicity , the weighting of each model corresponds to the number of systems resembling that model expected in a stellar population of M⊙. For full details about the binary stellar models and the code, we refer the reader to the bpass v2 release paper 2017PASA...34...58E.
@@
For the metallicity-dependent cosmic star formation rate history, CSFH, we adopt the model of 2006ApJ...638L..63L. We seed the population synthesis models according to this CSFH such that the volumetric birth rate of each system (bpass model) at each redshift and metallicity is proportional to the product of the star formation rate density SFRD and the model weighting (defined by the IMF and binary parameter distributions as described above).
@@
We adopt the standard bpass assumptions that core-collapse occurs when, at the end of the model, the total mass exceeds 1.5 M⊙, the CO core mass exceeds 1.38 M⊙ and the ONe core mass is non-zero. Remnant masses are a pre-calculated output of the models. They are determined by injecting 1051 erg of energy, calculating the mass in the outer layers which can be lifted to infinity by this energy injection, and taking the remaining mass as the remnant mass (for full details see 2004MNRAS.353...87E; 2017PASA...34...58E). We note that this energy injection is likely an overestimate for BH-forming events and means we have a large amount of mass potentially available for fallback accretion (this is addressed further in the discussion). We classify all core-collapse remnants with M3 M⊙ as neutron stars, and those heavier as black holes. We assume that all NS-forming events produce successful supernovae, and that BH-forming events do not. Pair-instability supernovae (PISNe) are deemed to occur if the final carbon-oxygen core mass exceeds 60 M⊙ and the final helium core mass is 133 M⊙. In this case no remnant is left behind (2002ApJ...567..532H). Pulsational pair instability supernovae (PPISNe) are accounted for by adopting the remnant mass prescription of 2019ApJ...887...53F for final carbon-oxygen masses between 38 and 60 M⊙, following 2023MNRAS.520.5724B. The true landscape of explodability and remnant formation is more complex, and our assumptions neglect islands of explodability above the mass regime where neutron stars are formed, but capture the broad behaviour predicted by many works.(e.g. 2009MNRAS.395.1409S; 2009ARA&A..47...63S; 2012ApJ...749...91F; 2020ApJ...890...51E; 2020MNRAS.492.2578S; 2020MNRAS.499.2803P; 2021ApJ...909..169K; 2022MNRAS.511..903P; 2025A&A...695A..71L; 2025A&A...695A.122U; 2025A&A...700A..20M). These assumptions also naturally reproduce observed (binary) remnant mass distributions (e.g. 2014ApJ...785...28K; 2023MNRAS.520.5724B; 2023A&A...676A..31D; 2025arXiv250818083T). Nevertheless, we acknowledge the possible impacts of these simplifying assumptions in Section 5.3.
@@
The remnant formation time for each stellar model is the seed time, plus the age of the model at core-collapse (although for these short-lived stars, this is negligible compared with cosmological timescales). In this way, we construct a metallicity-weighted supernova rate and BH formation rate as a function of redshift. Since all (confirmed) LFBOTs thus far have redshifts , we compare the mean rate of supernovae and BH-forming events over this redshift range, for different values of MBH,min. Specifically, we calculate the mean ratio of the supernova rate and the formation rate of BHs with mass MBH,min or greater. As the form of CSFH is uncertain, we repeat the process with two extreme variations of the CSFH. These are defined by shifting the 2006ApJ...638L..63L model by 0.2 dex in 12log(O/H), which captures the approximate envelope of possible CSFHs found by 2019MNRAS.488.5300C.
@@
The results are shown in Figure 1. Using the fiducial CSFH we find that the birth rate of BHs at with masses 38 M⊙ is 0.1% of the CCSN rate, matching LFBOTs. The progenitor temperatures, luminosities and radii pre-collapse are provided in Figure 7 of Appendix A. For a CSFH of the same form but peaking at a metallicity 0.2 dex lower in 12log(O/H), we find a higher BH mass threshold of 41 M⊙ is required, because there are more high-mass BHs being formed overall with respect to CCSNe. The reverse is true for a CSFH which peaks 0.2 dex higher, yielding a threshold of 33 M⊙. Therefore, varying the CSFH changes the absolute number of BHs being born (although not the general shape of the BH mass distribution, 2023ApJ...948..105V), and the number formed with respect to CCSNe.
@@ 3 Metallicity dependence and LFBOT host galaxies
$a$$a$footnotetext: IFU metallicity measurement, resolves the host giving a local (few hundred parsec scale) value.$b$$b$footnotetext: Metallicity derived from the host’s total stellar mass using equation 3 of 2004ApJ...613..898T. Uncertainties are solely propagated from uncertainties on the galaxy stellar mass, and do not include scatter in the relation.
@@
| AT 2018cow | 0.470.01 | 2020MNRAS.495..992La |
| ZTF18abvkwla | 0.37 | 2020ApJ...895...49H |
| CSS161010 | 0.15 | 2024ApJ...977..162G |
| ZTF20acigmel | 0.15 | 2021MNRAS.508.5138Pb |
| AT2020mrf | 0.27 | 2022ApJ...934..104Y |
| AT2023fhn | 0.080.02 | 2024AandA...691A.329C |
| AT2024wpp | 0.53 | 2025MNRAS.537.3298Pb |
@@
Given a black hole mass cut MBH,min which yields a BH formation rate matching the observed LFBOT rate, for an assumed CSFH, we can next ask: what is the (CSFH weighted) metallicity distribution of the progenitors of these black holes, and how does this compare with the host galaxies of LFBOTs? We collate the available information in the literature on LFBOT host metallicities in Table 1. Four of the seven have host-averaged metallicities derived from spectral energy distribution (SED) fitting. In one case (AT 2018cow) there is integral field unit data available, enabling a more local (resolution of a few hundred parsec, 2020MNRAS.495..992L) measurement of the metallicity. For the remaining two, no direct measurement has been published, so we obtain an approximate value through the galaxy mass-metallicity relation (2004ApJ...613..898T). We adopt Z by mass fraction for Solar metallicity, but note that this specific choice has no impact on the comparisons or analysis since both the observed and synthetic metallicities are scaled by the same factor.
@@
In Figure 2 we compare the number of black holes more massive than MBH,min born per M⊙ of star formation, weighted by the mean metallicity distribution for each of the three CSFHs adopted, with the measured distribution of LFBOT host or environmental metallicities. There is a strong preference for BHs with masses greater than 38 M⊙ to be born in stellar populations with metallicity less than 0.3 Z⊙. Below this threshold, the distributions are relatively constant: this reflects that the fact that (i) we are tailoring our mass cut to yield a BH formation rate that is 0.1% of the CCSN rate for each CSFH, and (ii) that the CCSN rate itself is not substantially varying with the CSFH variations. This is because NS formation efficiency, unlike BH formation efficiency, does not sharply decline at higher metallicities (e.g. 2025ApJ...979..209V).
@@
It is immediately apparent that LFBOT environments sample the ZZ⊙ region. The formation of BHs with MMBH,min across this range instead samples metallicities ZZ⊙. A KS-test between the LFBOT host values and the fiducial M⊙ distribution returns , formally rejecting the null hypothesis that the observed host values are consistent with the synthetic distribution. However, we note that the host metallicities are far from uniformly determined - with methods ranging from resolved IFU spectroscopy, probing the local environment within the host, to host-integrated metallicites to the mass-metallicity relation. This likely introduces errors/scatter which has not been accounted for, and host-integrated measurements will naturally be higher than the lowest metallicity environments within each host. There may also be pockets of low metallicity gas within and around star-forming galaxies, driving low-Z star-formation (e.g. 2015A&A...582A..78M; 2023MNRAS.520..879M). Nevertheless, the current information suggesting that LFBOTs can occur in environments with metallicities above the strong cut-off predicted is a challenge to the model. This is most notable for AT2018cow, whose metallicity measurement is integrated across the immediate (few hundred pc) environment (2020MNRAS.495..992L), but even in this case variations on smaller scales are possible.